# BabyLM’s First Words: Word Segmentation as a Phonological Probing Task

### v2

@@ 1 Introduction
Figure 1: Three strategies for unsupervised word segmentation using cues extracted from an autoregressive language model trained to predict phonemes.
@@ 2 Related Work and Motivations
Early “connectionist” language models were trained on sequences of letters or phonemes, often using developmentally plausible data in order to explore theories of word learning and phonology (Seidenberg and McClelland, 1989; Norris, 1994; Coltheart et al., 2001). Modern large language models (LLMs) are still probed for grammatical information, but standard benchmarks are generally based on higher-order structures: syntax and semantics rather than morphology and phonology. This due to LLM design being optimized for downstream tasks, not linguistic analysis. For instance, LLMs are typically trained on graphemic text using subword tokens. While this representation is practical for large-scale training, these tokens are not very cognitively plausible (Beinborn and Pinter, 2023), are less effective than character-based tokens for learning word structure (Bunzeck and Zarrieß, 2025) and cannot be used to explore representations of phonological units. Additionally, modern LLMs are inappropriate for theories of acquisition, due to the scales of data they are trained on (Warstadt et al., 2023).
@@ 2.2 Computational Models of Segmentation
Unlike in written text, where lexical units are separated by spaces and punctuation, spoken communication consists of continuous utterances with no clear demarcation of words (see e.g. Cole and Jakimik (1980)). Somehow, without a lexicon to consult, children are able to segment speech into words and phrasal units by the age of six months (Jusczyk, 1999). How children learn to segment words and bootstrap their lexicon is known in psycholinguistics as the word segmentation problem, and statistical learning experiments have established a wide variety of statistical cues which children may use to segment speech (Cutler and Carter, 1987; Gleitman et al., 1988; Jusczyk et al., 1993; Saffran et al., 1996b; Jusczyk et al., 1999; Suomi et al., 1997).
Particularly influential were the experiments of Saffran et al. (1996a), who established that 8-month-old children use distributional information to segment speech, specifically noting that low conditional probability between two adjacent syllables often indicated a word boundary. These experiments inspired the development of computational models proposing cognitively plausible learning mechanisms for word segmentation, most of which are based on the principle that units within words are far more predictable than units across word boundaries (Harris, 1955). Many models draw on Brent (1999), who used unigram statistics to segment speech, with later models using higher-order n-grams (Venkataraman, 2001), incorporating phonological constraints (Blanchard et al., 2010) or leveraging prior distributions over word frequencies and phonological shapes (Goldwater et al., 2009). Other models explicitly calculate several statistical cues at each potential word boundary and combine cues using a majority voting framework (Çöltekin and Nerbonne, 2014; Çöltekin, 2017; Goriely et al., 2023). Each cue provides a signal over the utterance (as illustrated in fig. 1) with peaks in each cue indicating a potential boundary.
Peaks in predictability can also be observed in neural language models. In the foundational work of Elman (1990), a simple recurrent network (SRN) is trained to predict letters in an unsegmented sequence (one of the first examples of autoregressive language modeling). Elman observes that the prediction-error increases at the onset of each new word, concluding that “there is information in the signal that could serve as a cue to the boundaries of linguistic units which must be learned”.
@@ 4 Experimental Setup
Table 1: The model size in number of (non-embedding) parameters and data size used for each suite of models. Languages are subsampled according to the token count for consistency, as word length varies across languages.
@@ 6 Discussion
In this work, we train BabyLMs on phonemic transcriptions of 31 languages in IPA CHILDES and explore the word segmentation task as a method for probing these models for phonological knowledge. Our results indicate that prediction-error and utterance boundary probability can be used as cues for unsupervised word segmentation. Our study is the first to use prediction-error extracted from LLMs for unsupervised word segmentation, extending previous work that explicitly calculated these cues using n-gram models (Çöltekin and Nerbonne, 2014; Çöltekin, 2017; Goriely et al., 2023). We also update previous neural models of word recognition (Elman, 1990; Christiansen et al., 1998) by using modern architectures and evaluating cross-lingually.
@@
Comparing models across languages is a challenge. Our study is the first cross-lingual word segmentation study to compare 31 languages, but we identify two confounding factors that inhibit cross-lingual comparison. Firstly, we find that the distribution of phonemes in word-final slots provides a prior not previously accounted for in studies that probed contextual embeddings for word boundary information. Secondly, we find that word length provides a prior for the unsupervised strategies, since randomly placing boundaries yields a higher F1 score when words are shorter, which has not previously been accounted for in cross-lingual word segmentation studies. Nevertheless, both the probes and the unsupervised strategies achieve significant scores for all 31 languages, indicating the importance of the distributional cue for learning to segment speech in any language. These findings also highlight the importance of accounting for frequency information as a prior when training probes or comparing models trained on different datasets.
@@ Appendix A Limitations
Using phonemic data for the word segmentation task is the typical framework for exploring relevant acquisition theories. However, the phonemic transcriptions in IPA CHILDES do have limitations. Having been generated using grapheme-to-phoneme (G2P) conversion, they may have been subject to conversion error, and the original transcriptions may also contain errors. The G2P process also removes natural variation in speech, such as accents and allophonic variation. The symbolic nature of phonemes may also be an unrealistic starting point for acquisition; it is unclear if infants have access to phonetic categories at this stage of acquisition (Feldman et al., 2021; McMurray, 2022). When simulating acquisition, a better framework may be to learn to segment directly from audio, as explored in the Zero Resource Speech Challenge (Nguyen et al., 2020; Dunbar et al., 2021). We do not claim that our phoneme LMs simulate acquisition. Instead, they serve as a useful method for studying phonology at the phoneme level, and examining the distributional properties of speech that could be available to language-learning infants.

### v3

@@ 1 Introduction
Figure 1: Three strategies for unsupervised word segmentation using cues extracted from an auto-regressive language model trained to predict phonemes.
@@ 2 Related Work and Motivations
Early “connectionist” language models were trained on sequences of letters or phonemes, often using developmentally plausible data in order to explore theories of word learning and phonology (Seidenberg and McClelland, 1989; Norris, 1994; Coltheart et al., 2001). Modern large language models (LLMs) are still probed for grammatical information, but standard benchmarks are generally based on higher-order structures: syntax and semantics rather than morphology and phonology. This is due to LLM design being optimized for downstream tasks, not linguistic analysis. For instance, LLMs are typically trained on graphemic text using subword tokens. While this representation is practical for large-scale training, these tokens are not very cognitively plausible (Beinborn and Pinter, 2023), are less effective than character-based tokens for learning word structure (Bunzeck and Zarrieß, 2025) and cannot be used to explore representations of phonological units. Additionally, modern LLMs are inappropriate for theories of acquisition, due to the scales of data they are trained on (Warstadt et al., 2023).
@@ 2.2 Computational Models of Segmentation
Unlike in written text, where lexical units are separated by spaces and punctuation, spoken communication consists of continuous utterances with no clear demarcation of words (see, e.g. Cole and Jakimik, 1980). Somehow, without a lexicon to consult, children are able to segment speech into words and phrasal units by the age of six months (Jusczyk, 1999). How children learn to segment words and bootstrap their lexicon is known in psycholinguistics as the word segmentation problem, and statistical learning experiments have established a wide variety of statistical cues which children may use to segment speech (Cutler and Carter, 1987; Gleitman et al., 1988; Jusczyk et al., 1993; Saffran et al., 1996b; Jusczyk et al., 1999a; Suomi et al., 1997).
Particularly influential were the experiments of Saffran et al. (1996a), who established that 8-month-old children use distributional information to segment speech, specifically noting that low conditional probability between two adjacent syllables often indicated a word boundary. These experiments inspired the development of computational models proposing cognitively plausible learning mechanisms for word segmentation, most of which are based on the principle that units within words are far more predictable than units across word boundaries (Harris, 1955). Many models draw on Brent (1999), who use unigram statistics to segment speech, with later models using higher-order n-grams (Venkataraman, 2001), incorporating phonological constraints (Blanchard et al., 2010) or leveraging prior distributions over word frequencies and phonological shapes (Goldwater et al., 2009). Other models explicitly calculate several statistical cues at each potential word boundary and combine cues using a majority voting framework (Çöltekin and Nerbonne, 2014; Çöltekin, 2017; Goriely et al., 2023). Each cue provides a signal over the utterance (as illustrated in fig. 1) with peaks in each cue indicating a potential boundary.
Peaks in predictability can also be observed in neural language models. In the foundational work of Elman (1990), a simple recurrent network (SRN) is trained to predict letters in an unsegmented sequence (one of the first examples of auto-regressive language modeling). Elman observes that the prediction-error increases at the onset of each new word, concluding that “there is information in the signal that could serve as a cue to the boundaries of linguistic units which must be learned”.
@@
While our experimental setup draws on previous computational work in word segmentation, we do not claim that our phoneme-level language models simulate child language acquisition (see section 6). Rather, we use the segmentation task — with phoneme-level input — as a diagnostic tool that allows us to characterize the cross-linguistic distributional structure of speech sounds and test whether language models naturally group cluster sequences into units that coincide with our notion of word-hood. Although our findings may support aspects of statistical learning theories, we acknowledge the limitations of using phoneme-based representations in appendix A.
@@ 4 Experimental Setup
Table 1: The model size in number of (non-embedding) parameters and data size used for each suite of models. Languages are sub-sampled according to the token count for consistency, as word length varies across languages.
@@ 6 Discussion
In this work, we train BabyLMs on phonemic transcriptions of 31 languages in IPA CHILDES and explore the word segmentation task as a method for probing these models for phonological knowledge. Our results indicate that prediction-error and utterance boundary probability can be used as cues for unsupervised word segmentation. Our study is the first to use prediction-error extracted from LLMs for unsupervised word segmentation, extending previous work that explicitly calculated these cues using n-gram models (Çöltekin and Nerbonne, 2014; Çöltekin, 2017; Goriely et al., 2023). We also update previous neural models of word recognition (Elman, 1990; Christiansen et al., 1998) by using modern architectures and evaluating cross-lingually. We now turn to the broader implications of our findings.
Statistical learning.
Viewing our models as statistical learners, we find that no single cue or strategy consistently yields the best segmentation performance across different model sizes and languages. This is perhaps unsurprising, as many of the cues are highly interrelated (for example, entropy and surprisal often correlate) and all segmentation strategies are grounded in the same underlying principle: identifying boundaries at points of high prediction uncertainty. It is this general principle, rather than any specific cue or strategy, that proves sufficient for segmenting utterances into word-like units. Nevertheless, most cues and strategies perform reasonably well on their own. Previous segmentation models have explored combining multiple distributional cues through unsupervised majority voting (Çöltekin, 2017; Goriely et al., 2023), an approach that could be fruitfully applied to the cues investigated here in future work.
Cross-lingual comparison.
@@
Comparing models across languages is a challenge. Our study is the first cross-lingual study using the word segmentation task to compare 31 languages, but we identify two confounding factors that inhibit cross-lingual comparison. Firstly, we find that the distribution of phonemes in word-final slots provides a prior not previously accounted for in studies that probed contextual embeddings for word boundary information. Secondly, we find that word length provides a prior for the unsupervised strategies, since randomly placing boundaries yields a higher F1 score when words are shorter, which has not previously been accounted for in cross-lingual word segmentation studies. Nevertheless, both the probes and the unsupervised strategies achieve significant scores for all 31 languages, indicating the importance of the distributional cue for learning to segment speech in any language. These findings also highlight the importance of accounting for frequency information as a prior when training probes or comparing models trained on different datasets.
Simulating acquisition.
Our results focus on the performance of our models at the end of training, whereas past work has compared the learning dynamics of phoneme-based models to developmental patterns observed in human acquisition (Kirov and Cotterell, 2018). Although our findings indicate the utility of the distributional cue for identifying word-like units, we do not claim that our models simulate language acquisition. In particular, given recent advances in models that operate directly on raw audio, the use of phoneme-level representations may be insufficient for capturing the full complexity of language learning, as discussed in appendix A.
Rather, we use this framework to investigate the distributional patterns of phonemes across languages and whether language models trained to predict upcoming phonemes implicitly track meaningful sub-sequences that align with words. While many computational models of word segmentation treat segmentation as a necessary precursor for language understanding, this assumption has been questioned. For example, Baayen et al. (2016) show that a tri-phone model, operating on unsegmented utterances can make predictions consistent with infants’ sensitivity to linguistic structure. Likewise, recent phoneme-level language models perform well on both linguistic benchmarks and downstream tasks without explicit segmentation (Goriely et al., 2024) — although our results suggest that some degree of implicit segmentation may be occurring to enhance these models’ predictive performance.
Word boundaries as gold labels.
@@ Appendix A Limitations
Using phonemic data for the word segmentation task is the typical framework for exploring relevant acquisition theories. However, the phonemic transcriptions in IPA CHILDES do have limitations. Having been generated using grapheme-to-phoneme (G2P) conversion, they may have been subject to conversion error, and the original transcriptions may also contain errors. The G2P process also removes natural variation in speech, such as accents and allophonic variation. The symbolic nature of phonemes may also be an unrealistic starting point for acquisition; it is unclear if infants have access to phonetic categories at this stage of acquisition (Feldman et al., 2021; McMurray, 2022). Researchers who advocate for using language models as cognitive models argue that the training data should be as developmentally plausible as possible (Dupoux, 2018; Warstadt and Bowman, 2022), and that phonemes may be as implausible as text for simulating early acquisition (Lavechin et al., 2023).
From this perspective, a more appropriate framework is to learn segmentation directly from raw audio, as pursued in the Zero Resource Speech Challenge (Nguyen et al., 2020; Dunbar et al., 2021). Audio-based models naturally incorporate prosodic cues, which play a key role in language acquisition (Cutler and Carter, 1987; Jusczyk et al., 1993, 1999b). Unsupervised models have demonstrated the ability to perform statistical learning directly from raw speech (Lavechin et al., 2022; de Seyssel et al., 2023), and have found that the resulting units tend to be shorter than phonemes, consistent with early perceptual categories (Schatz et al., 2021). While such models show promising signs of early phonetic learning and perform well on word-level tasks, they currently require significantly more data to match the performance of text-based models (Lavechin et al., 2023). Moreover, training on curated audiobook datasets gives these models a considerable advantage over learning from noisier, long-form audio that better resembles real-world input—but ongoing work is making such realistic simulations increasingly viable (Lavechin et al., 2024).